4. Determine the quadratic function,f(x)=ax˛+bx+c , whose vertex is (1, -2) and passes through the point (-1, -22). Please show all of your work.
I appreciate any help! Thanks!

Answer by stanbon(75887)

Determine the quadratic function,f(x)=ax˛+bx+c , whose vertex is (1, -2) and passes through the point (-1, -22).----------------
---
y = a(x-h)^2+k
---
(h,k) = (1,-2)
---
y = a(x-1)^2-2
---
Use (-1,-22) to solve for "a":
-22 = a(-1-1)^2-2
-22 = a(2)
a = -11
----
Equation:
y = -11(x-1)^2-2
----
y = -11(x^2-2x+1)-2
---
y = -11x^2+22x-13
=====================
Cheers,
Stan H.
============